# Tidal perturbations of an extreme mass ratio inspiral around a Kerr black hole

### v1

@@ 3.1 Hertz potential
Here $\tilde{v}$ denotes a slowly varying time variable, in the sense that the tidal field is assumed to vary on timescales much longer than the characteristic dynamical timescale of the black hole. As a consequence, time derivatives are negligible compared to spatial derivatives [Binnington:2009bb, LeTiec:2020bos]. Accordingly, the Teukolsky equation is solved in the static ($\omega=0$) sector, and the coefficients $c_{2m}$ depend on $\tilde{v}$ only parametrically.
@@
where $\mathbf{F}(a,b,c;z)$ denotes the regularized hypergeometric function, and $x$ and $\gamma$ are defined in Eq. (2.24). This expression coincides with the result obtained in the Kinnersley tetrad [LeTiec:2020bos], up to the expected rescaling by the square of the factor $\Delta/[2(r^{2}+a^{2})]$ relating the Kinnersley and HH tetrads, in agreement with Eq. (2.11).
@@
From the perspective of the external spacetime, the black hole can be treated as a point particle moving along a worldline $\mathscr{L}$, which is endowed with an orthonormal tetrad $(u^{\alpha},e^{\alpha}_{a})$. Here $u^{\alpha}$ is the black-hole four-velocity and $e^{\alpha}_{a}$ is a spatial triad parallel transported along $\mathscr{L}$. This tetrad defines the local asymptotic rest frame of the black hole within the external geometry. Projecting the external Weyl tensor onto this frame yields the components [Zhang:1986cpa, Poisson:2009qj, Binnington:2009bb, Chatziioannou:2012gq, Poisson:2014gka]
@@
These frame components can in turn be recast in terms of two symmetric, trace-free tensors $\mathscr{E}_{ab}$ and $\mathscr{B}_{ab}$, representing the electric-type and magnetic-type quadrupole moments of the tidal field, respectively [Poisson:2009qj]. They are defined as
@@
As an alternative to the standard asymptotic-matching approach of Refs. [Poisson:2009qj, Binnington:2009bb, Poisson:2004cw, Chatziioannou:2012gq], we determine the coefficients $c_{2m}$ by requiring that the metric reconstructed from $\Psi^{\prime}_{\rm H}$ reproduces the known quadrupolar tidal perturbation computed in Ref. [Binnington:2009bb] in the Schwarzschild limit $a=0$. The two procedures are equivalent and consistent with each other once the relation between the Weyl scalar $\psi_{0}$ and the Hertz potential $\Psi^{\prime}_{\rm H}$ is taken into account (see, e.g., Eq. (1.23a) of Ref. [Berens:2024czo]). The details of our procedure, along with the final expressions for $c_{2m}$, are presented at the end of Sec. 4.
@@ 4 Perturbation to the Kerr metric: results
The final step, as indicated in Eq. (3.2), is to take the real part of this expression, yielding $h_{\mu\nu}^{\rm ORG}\equiv 2\,\mathrm{Re}\!\left[\hat{h}_{\mu\nu}^{\rm ORG}\right]$. The explicit expressions for each $m$-mode are given below. To simplify the notation, we henceforth omit the superscript “ORG”. We present the full expressions explicitly in order to facilitate direct reuse in analytic and numerical applications.
@@ 4.3 $m=\pm 2$ modes
The complete perturbation $h_{\mu\nu}^{\rm ORG}$, obtained by summing the five contributions above and taking the real part of the result, satisfies the ORG conditions. One can further verify that, in the zero-spin limit $a=0$, it reduces to the known quadrupolar tidal perturbation of the Schwarzschild metric derived in Ref. [Binnington:2009bb]. Moreover, by expanding the reconstructed metric to first order in the spin parameter $a$, the associated Weyl scalars $\psi_{0}$ and $\psi_{4}$ agree with those obtained in Refs. [Poisson:2014gka, Landry:2015zfa].
Finally, we have explicitly verified that each mode of the metric perturbation $\hat{h}_{\mu\nu}^{(m)\,\rm ORG}$ satisfies the linearized Einstein equations. For completeness, the full expressions for the metric perturbation components are provided in electronic form as a Mathematica Notebook in the supplementary material to this paper.
By performing a suitable gauge transformation from the ORG to the IRG and rotating from the HH to the Kinnersley tetrad, the tidally perturbed Kerr metric constructed here can be related to those obtained in Refs. [LeTiec:2020bos, Yunes:2005ve]. We also note that explicit reconstruction formulas in both the IRG and ORG within the Kinnersley tetrad were recently presented in Ref. [Berens:2024czo]. [^2] See Ref. [Berens:2025kkm] for an extension of this formalism to Kerr black holes embedded in a spacetime with a nonzero cosmological constant. Those expressions, however, were neither specialized to tidal perturbations nor given in the fully explicit analytic form derived here, which facilitates direct use in astrophysical applications (see Secs. 5 and 6).
@@ 4.4 Fixing the coefficients $c_{2m}$
As anticipated in Sec. 3.1, rather than following the standard asymptotic-matching procedure employed in Refs. [Poisson:2009qj, Binnington:2009bb, Poisson:2004cw, Chatziioannou:2012gq], we take the zero-spin limit ($a=0$) of the metric perturbation (3.2) and require that it reproduces the quadrupolar order of the tidally deformed metric derived in Ref. [Binnington:2009bb] for a Schwarzschild black hole. This fixes the coefficients $c_{2m}$ in terms of the external tidal moments.
@@
As anticipated in Sec. 3, once the relation between the Weyl scalar $\psi_{0}$ and the Hertz potential $\Psi^{\prime}_{\rm H}$ is taken into account, these coefficients are consistent with those obtained in Refs. [Poisson:2004cw, Chatziioannou:2012gq].
Finally, for later convenience, we introduce the scalar tidal potentials [Poisson:2014gka]
@@ 5 Secular dynamics in tidally deformed Kerr spacetime
A compelling and not yet fully explored question concerns how self-gravitating systems respond dynamically to external tidal fields. While tidal Love numbers have been extensively studied for compact objects in a wide range of theoretical frameworks [Chakraborty:2025wvs, Chia:2020yla, LeTiec:2020bos, LeTiec:2020spy, Pani:2015hfa, Gurlebeck:2015xpa, Poisson:2014gka, Kol:2011vg, Damour:2009va, Landry:2015zfa, Pani:2015nua, Binnington:2009bb, Chakrabarti:2013lua, Damour:2009vw, Chatziioannou:2020pqz, Flanagan:2007ix, Hinderer:2007mb, Cardoso:2017cfl, Pereniguez:2025jxq, Charalambous:2021mea, Perry:2023wmm], comparatively little attention has been devoted to the dynamics of matter orbiting a tidally distorted compact object.
Notable exceptions include studies of EMRI systems subject to external perturbations, which have shown modifications to orbital evolution and the possible onset of tidal resonances [Yang:2017aht, Yang:2019iqa, Bonga:2019ycj, Gupta:2021cno, Gupta:2022fbe].
@@
In this regime, the dynamics is naturally characterized by a separation of timescales between the fast orbital motion and the much slower evolution induced by weak external perturbations. This hierarchy provides the basis for a description in terms of quantities averaged over the orbital period [Mino:2003yg, Hinderer:2008dm].
@@ 5.1 Secular Hamiltonian
We begin by examining the case of a test particle of mass $m$ moving on a timelike geodesic in the tidally deformed spacetime with metric $g_{\mu\nu}$. To describe the dynamics of this particle, we use the geodesic equation, including corrections to the Kerr geometry up to linear order in the metric perturbation $h_{\mu\nu}$. This approximation is valid provided that self-force effects can be neglected, namely in the limit $m/M\ll 1$, where $M$ denotes the mass of the Kerr black hole [Fujita:2016igj]. We also require $M/\mathscr{R}\ll 1$, where $\mathscr{R}$ is the length scale associated with the external tidal field, and we retain contributions only up to leading quadrupolar order, $\mathscr{O}\!\left((M/\mathscr{R})^{2}\right)$.
@@
For bound geodesic motion in Kerr spacetime, the radial, polar, and azimuthal motions are characterized by three independent fundamental frequencies. It is therefore convenient to parametrize the orbit using three phase variables $(q_{r},q_{z},q_{\psi})$, which label the oscillations associated with the radial, polar, and azimuthal motions, respectively, and evolve linearly with respect to a suitable parameter along the geodesic [Hinderer:2008dm, vandeMeent:2019cam, Schmidt:2002qk, Drasco:2003ky].
For a generic, non-resonant orbit, the secular average of a quantity $\mathscr{A}$ is defined as [Drasco:2005is]
@@
As already mentioned, at the order considered here, the secularly averaged geodesic $\Gamma$ coincides with the physical orbit $\Gamma^{\prime}$. More precisely, differences arise only at higher order, for example when including first-order corrections in the mass ratio $m/M$ or second-order contributions in the metric perturbation $h_{\mu\nu}$ [Fujita:2016igj, Yang:2017aht, Camilloni:2023rra].
Moreover, to first order in the tidal perturbation, the secularly averaged energy $E$ and angular momentum $L$, both defined per unit rest mass, remain conserved along the secular orbit $\Gamma$ [Fujita:2016igj, Camilloni:2023rra].
@@
$\langle h_{vr}\rangle~=$ $~\langle h_{rr}\rangle~=\langle h_{rz}\rangle=\langle h_{r\phi}\rangle=0~,$ (5.5g)
where, using Eq. (4.12), the electric tidal scalar $\mathscr{E}^{q}$ is written explicitly as
@@
and its secular average on the equatorial plane is given by $\langle\mathscr{E}^{q}\rangle\equiv(\mathscr{E}_{11}+\mathscr{E}_{22})/2$. Equivalently, one obtains an expression analogous to (5.6) for the magnetic tidal scalar $\mathscr{B}^{q}$, and its secular average on the equatorial plane is given by $\langle\mathscr{B}^{q}\rangle\equiv(\mathscr{B}_{11}+\mathscr{B}_{22})/2$.
@@
is a dimensionless perturbative parameter encoding information about the tidal environment [Yang:2017aht, Camilloni:2023rra, Grilli:2024fds]. We note that, although some components of the metric perturbation depend on the magnetic tidal scalar $\mathscr{B}^{\rm q}$, even after the averaging procedure, these components do not contribute to the Hamiltonian. As a result, any dependence on $\langle\mathscr{B}^{\rm q}\rangle$ drops out.
@@
In the zero-spin limit $a=0$, Eqs. (5.9) and (5.10) reproduce the results obtained in Refs. [Yang:2017aht, Camilloni:2023rra] for a tidally deformed Schwarzschild spacetime.
@@ 6 Tidal deformation of ISCO and LR
The ISCO is not only a mathematically well-defined geodesic of the Kerr geometry [Bardeen:1972fi], but also an object of direct astrophysical relevance. The binding energy at the ISCO sets the radiative efficiency of thin accretion disks, and measurements of disk spectra have therefore been used to infer black-hole spins through continuum-fitting methods [Narayan:2005ie, Zhang:1997dy]. Similarly, the $\mathrm{K}\alpha$ iron line originates in the innermost regions of the accretion disk, and its profile is sensitive to the location of the disk inner edge, which is often identified with the ISCO [Fabian:1989ej, Fabian:2000nu].
From a gravitational-wave perspective, the ISCO marks the end of the adiabatic inspiral and the onset of the plunge phase. In EMRI systems, even small tidal shifts in the location of the ISCO or in the associated orbital constants can accumulate over many cycles and lead to potentially detectable dephasings in the gravitational-wave signal [Ori:2000zn].
The LR is likewise of strong physical and observational relevance. Early studies showed that the LR controls the optical appearance of a collapsing star and the exponential fade-out of its observed luminosity [Ames:1968apj]. More recently, the LR has been recognized as playing a pivotal role in a variety of observational and theoretical contexts, including the determination of the photon-ring structure and, consequently, of the black-hole shadow [Falcke:1999pj]. The angular frequency and instability timescale (i.e., Lyapunov exponent) of the LR are directly related to the real and imaginary parts of black-hole quasinormal modes in the eikonal limit [Goebel:1972apj, Ferrari:1984zz, Mashhoon:1985cya, Berti:2005eb, Cardoso:2008bp, Cardoso:2021qqu, Yang:2012he]. Furthermore, the LR also determines the critical impact parameter relevant for high-energy scattering and threshold phenomena in numerical collision experiments [Pretorius:2007jn].
Computing tidal-induced secular shifts of LR quantities, such as the radius, orbital frequency, impact parameter, and instability rate, is therefore important, as such shifts could leave measurable imprints on black-hole shadows, ringdown signals, and critical scattering dynamics [Cardoso:2021qqu].
@@ 6.1 ISCO shifts
We begin by computing the shift induced by the tidal deformation on the location and properties of the ISCO of a Kerr black hole. We recall that, in the equatorial plane of the Kerr spacetime, there exist two ISCOs corresponding to prograde and retrograde orbits, respectively. These orbits can be determined by imposing the following three conditions [Isoyama:2014mja]
@@
where $\sigma=\pm 1$ labels prograde ($\sigma=+1$) and retrograde ($\sigma=-1$) orbits, respectively. Analogously, one can compute the orbital frequency $\Omega\equiv\langle u^{\psi}/u^{v}\rangle$ and the redshift invariant $U\equiv\langle u^{v}\rangle$ [Detweiler:2008ft, Sago:2008id, Shah:2012gu, vandeMeent:2015lxa], [^4] For circular equatorial orbits, these quantities coincide with $\langle u^{\phi}/u^{t}\rangle$ and $\langle u^{t}\rangle$ in Boyer-Lindquist coordinates, since $(v,\psi)$ differ from $(t,\phi)$ only by functions of the radial coordinate $r$ (cfr. Eq. (2.2)).
@@
By solving Eqs. (6.1) at zeroth order in the tidal parameter $\eta$, one recovers the well-known expressions for the radial coordinate, the specific energy, and the specific angular momentum of the unperturbed ISCOs of a Kerr black hole, in agreement with the seminal results of Refs. [Chandrasekhar:1985kt, Bardeen:1972fi]:
@@
These expressions reveal a strong dependence of the tidal ISCO shifts on both the black-hole spin and the relative orientation of the orbit. In Fig. 1 we illustrate the dependence of these quantities on the spin of the Kerr black hole. We comment further on these results below. In the non-spinning limit $a=0$, all tidal corrections smoothly reduce to those computed for a tidally deformed Schwarzschild black hole in Ref. [Camilloni:2023rra].

### v2

@@ 3.1 Hertz potential
Here $\tilde{v}$ denotes a slowly varying time variable, in the sense that the tidal field is assumed to vary on timescales much longer than the characteristic dynamical timescale of the black hole. As a consequence, time derivatives are negligible compared to spatial derivatives [23, 21]. Accordingly, the Teukolsky equation is solved in the static ($\omega=0$) sector, and the coefficients $c_{2m}$ depend on $\tilde{v}$ only parametrically.
@@
where $\mathbf{F}(a,b,c;z)$ denotes the regularized hypergeometric function, and $x$ and $\gamma$ are defined in Eq. (2.24). It is worth noting that $R_{2m}(r)$ is an algebraic function of the Boyer–Lindquist radius $r$; indeed, notice that $\mathbf{F}(-4,1,-1+2im\gamma;-x)$ reduces to a polynomial. The Boyer–Lindquist radial coordinate is physically well-defined, since its level sets are totally-geodesic surfaces. This fact underlies the vanishing of black-hole tidal Love numbers in general relativity, and it extends to the general harmonic case $\ell\neq 2$, where $R_{\ell m}(r)$ reads
$R_{\ell m}(r)\sim\frac{1}{(r^{2}+a^{2})^{2}}\mathbf{F}(-2-\ell,-1+\ell,-1+2im\gamma;-x)\,,\qquad|m|\leq\ell,$ (3.10)
where $\mathbf{F}(-2-\ell,-1+\ell,-1+2im\gamma;-x)$ is again a polynomial. The expression in Eq. (3.9) coincides with the result obtained in the Kinnersley tetrad [21], up to the expected rescaling by the square of the factor $\Delta/[2(r^{2}+a^{2})]$ relating the Kinnersley and HH tetrads, in agreement with Eq. (2.11).
@@
From the perspective of the external spacetime, the black hole can be treated as a point particle moving along a worldline $\mathscr{L}$, which is endowed with an orthonormal tetrad $(u^{\alpha},e^{\alpha}_{a})$. Here $u^{\alpha}$ is the black-hole four-velocity and $e^{\alpha}_{a}$ is a spatial triad parallel transported along $\mathscr{L}$. This tetrad defines the local asymptotic rest frame of the black hole within the external geometry. Projecting the external Weyl tensor onto this frame yields the components [47, 37, 23, 13, 38]
@@
These frame components can in turn be recast in terms of two symmetric, trace-free tensors $\mathscr{E}_{ab}$ and $\mathscr{B}_{ab}$, representing the electric-type and magnetic-type quadrupole moments of the tidal field, respectively [37]. They are defined as
@@
As an alternative to the standard asymptotic-matching approach of Refs. [37, 23, 9, 13], we determine the coefficients $c_{2m}$ by requiring that the metric reconstructed from $\Psi^{\prime}_{\rm H}$ reproduces the known quadrupolar tidal perturbation computed in Ref. [23] in the Schwarzschild limit $a=0$. The two procedures are equivalent and consistent with each other once the relation between the Weyl scalar $\psi_{0}$ and the Hertz potential $\Psi^{\prime}_{\rm H}$ is taken into account (see, e.g., Eq. (1.23a) of Ref. [42]). The details of our procedure, along with the final expressions for $c_{2m}$, are presented at the end of Sec. 4.
@@ 4 Perturbation to the Kerr metric: results
The final step, as indicated in Eq. (3.2), is to take the real part of this expression, yielding $h_{\mu\nu}^{\rm ORG}\equiv 2\,\mathrm{Re}\!\left[\hat{h}_{\mu\nu}^{\rm ORG}\right]$. The explicit expressions for each $m$-mode are given below. To simplify the notation, we henceforth omit the superscript “ORG”. We present the full expressions explicitly in order to facilitate direct reuse in analytic and numerical applications. To this purpose, in the supplementary material, we also provide a Mathematica notebook with the components of the reconstructed metric.
@@ 4.3 $m=\pm 2$ modes
The complete perturbation $h_{\mu\nu}^{\rm ORG}$, obtained by summing the five contributions above and taking the real part of the result, satisfies the ORG conditions. One can further verify that, in the zero-spin limit $a=0$, it reduces to the known quadrupolar tidal perturbation of the Schwarzschild metric derived in Ref. [23]. Moreover, by expanding the reconstructed metric to first order in the spin parameter $a$, the associated Weyl scalars $\psi_{0}$ and $\psi_{4}$ agree with those obtained in Refs. [38, 48].
Finally, we have explicitly verified that each mode of the metric perturbation $\hat{h}_{\mu\nu}^{(m)\,\rm ORG}$ satisfies the linearized Einstein equations. For completeness, we recall that the full expression for the components of the metric perturbation is provided in electronic form as a Mathematica notebook in the supplementary material to this paper.
By performing a suitable gauge transformation from the ORG to the IRG and rotating from the HH to the Kinnersley tetrad, the tidally perturbed Kerr metric constructed here can be related to those obtained in Refs. [21, 22]. We also note that explicit reconstruction formulas in both the IRG and ORG within the Kinnersley tetrad were recently presented in Ref. [42]. [^3] See Ref. [49] for an extension of this formalism to Kerr black holes embedded in a spacetime with a nonzero cosmological constant. Those expressions, however, were neither specialized to tidal perturbations nor given in the fully explicit analytic form derived here, which facilitates direct use in astrophysical applications (see Secs. 5 and 6).
@@ 4.4 Fixing the coefficients $c_{2m}$
As anticipated in Sec. 3.1, rather than following the standard asymptotic-matching procedure employed in Refs. [37, 23, 9, 13], we take the zero-spin limit ($a=0$) of the metric perturbation (3.2) and require that it reproduces the quadrupolar order of the tidally deformed metric derived in Ref. [23] for a Schwarzschild black hole. This fixes the coefficients $c_{2m}$ in terms of the external tidal moments.
@@
As anticipated in Sec. 3, once the relation between the Weyl scalar $\psi_{0}$ and the Hertz potential $\Psi^{\prime}_{\rm H}$ is taken into account, these coefficients are consistent with those obtained in Refs. [9, 13].
Finally, for later convenience, we introduce the scalar tidal potentials [38]
@@ 5 Secular dynamics in tidally deformed Kerr spacetime
A compelling and not yet fully explored question concerns how self-gravitating systems respond dynamically to external tidal fields. While tidal Love numbers have been extensively studied for compact objects in a wide range of theoretical frameworks [50, 25, 21, 51, 39, 52, 38, 53, 54, 48, 55, 23, 56, 57, 34, 31, 32, 35, 58, 59, 60], comparatively little attention has been devoted to the dynamics of matter orbiting a tidally distorted compact object.
Notable exceptions include studies of EMRI systems subject to external perturbations, which have shown modifications to orbital evolution and the possible onset of tidal resonances [61, 62, 63, 64, 65].
@@
In this regime, the dynamics is naturally characterized by a separation of timescales between the fast orbital motion and the much slower evolution induced by weak external perturbations. This hierarchy provides the basis for a description in terms of quantities averaged over the orbital period [66, 67].
@@ 5.1 Secular Hamiltonian
We begin by examining the case of a test particle of mass $m$ moving on a timelike geodesic in the tidally deformed spacetime with metric $g_{\mu\nu}$. To describe the dynamics of this particle, we use the geodesic equation, including corrections to the Kerr geometry up to linear order in the metric perturbation $h_{\mu\nu}$. This approximation is valid provided that self-force effects can be neglected, namely in the limit $m/M\ll 1$, where $M$ denotes the mass of the Kerr black hole [68]. We also require $M/\mathscr{R}\ll 1$, where $\mathscr{R}$ is the length scale associated with the external tidal field, and we retain contributions only up to leading quadrupolar order, $\mathscr{O}\!\left((M/\mathscr{R})^{2}\right)$.
@@
For bound geodesic motion in Kerr spacetime, the radial, polar, and azimuthal motions are characterized by three independent fundamental frequencies. It is therefore convenient to parametrize the orbit using three phase variables $(q_{r},q_{z},q_{\psi})$, which label the oscillations associated with the radial, polar, and azimuthal motions, respectively, and evolve linearly with respect to a suitable parameter along the geodesic [67, 69, 70, 71].
For a generic, non-resonant orbit, the secular average of a quantity $\mathscr{A}$ is defined as [72]
@@
As already mentioned, at the order considered here, the secularly averaged geodesic $\Gamma$ coincides with the physical orbit $\Gamma^{\prime}$. More precisely, differences arise only at higher order, for example when including first-order corrections in the mass ratio $m/M$ or second-order contributions in the metric perturbation $h_{\mu\nu}$ [68, 61, 73].
Moreover, to first order in the tidal perturbation, the secularly averaged energy $E$ and angular momentum $L$, both defined per unit rest mass, remain conserved along the secular orbit $\Gamma$ [68, 73].
@@
$\langle h_{vr}\rangle~=$ $~\langle h_{rr}\rangle~=\langle h_{rz}\rangle=\langle h_{r\psi}\rangle=0~,$ (5.5g)
where, using Eq. (4.12), the electric tidal scalar $\mathscr{E}^{\rm q}$ is written explicitly as
@@
and its secular average on the equatorial plane is given by $\langle\mathscr{E}^{\rm q}\rangle\equiv(\mathscr{E}_{11}+\mathscr{E}_{22})/2$. Equivalently, one obtains an expression analogous to (5.6) for the magnetic tidal scalar $\mathscr{B}^{\rm q}$, and its secular average on the equatorial plane is given by $\langle\mathscr{B}^{\rm q}\rangle\equiv(\mathscr{B}_{11}+\mathscr{B}_{22})/2$.
@@
is a dimensionless perturbative parameter encoding information about the tidal environment [61, 73, 74]. We note that, although some components of the metric perturbation depend on the magnetic tidal scalar $\mathscr{B}^{\rm q}$, even after the averaging procedure, these components do not contribute to the Hamiltonian. As a result, any dependence on $\langle\mathscr{B}^{\rm q}\rangle$ drops out.
@@
In the zero-spin limit $a=0$, Eqs. (5.9) and (5.10) reproduce the results obtained in Refs. [61, 73] for a tidally deformed Schwarzschild spacetime.
@@ 6 Tidal deformation of ISCO and LR
The ISCO is not only a mathematically well-defined geodesic of the Kerr geometry [75], but also an object of direct astrophysical relevance. The binding energy at the ISCO sets the radiative efficiency of thin accretion disks, and measurements of disk spectra have therefore been used to infer black-hole spins through continuum-fitting methods [76, 77]. Similarly, the $\mathrm{K}\alpha$ iron line originates in the innermost regions of the accretion disk, and its profile is sensitive to the location of the disk’s inner edge, which is often identified with the ISCO [78, 79].
From a gravitational-wave perspective, the ISCO marks the end of the adiabatic inspiral and the onset of the plunge phase. In EMRI systems, even small tidal shifts in the location of the ISCO or in the associated orbital constants can accumulate over many cycles and lead to potentially detectable dephasings in the gravitational-wave signal [80].
The LR is likewise of strong physical and observational relevance. Early studies showed that the LR controls the optical appearance of a collapsing star and the exponential fade-out of its observed luminosity [81]. More recently, the LR has been recognized as playing a pivotal role in a variety of observational and theoretical contexts, including the determination of the photon-ring structure and, consequently, of the black-hole shadow [82]. The angular frequency and instability timescale (i.e., Lyapunov exponent) of the LR are directly related to the real and imaginary parts of black-hole quasinormal modes in the eikonal limit [83, 84, 85, 86, 87, 36, 88]. Furthermore, the LR also determines the critical impact parameter relevant for high-energy scattering and threshold phenomena in numerical collision experiments [89].
Computing tidal-induced secular shifts of LR quantities, such as the radius, orbital frequency, impact parameter, and instability rate, is therefore important, as such shifts could leave measurable imprints on black-hole shadows, ringdown signals, and critical scattering dynamics [36].
@@ 6.1 ISCO shifts
We begin by computing the shift induced by the tidal deformation on the location and properties of the ISCO of a Kerr black hole. We recall that, in the equatorial plane of the Kerr spacetime, there exist two ISCOs corresponding to prograde and retrograde orbits, respectively. These orbits can be determined by imposing the following three conditions [90]
@@
where $\sigma=\pm 1$ labels prograde ($\sigma=+1$) and retrograde ($\sigma=-1$) orbits, respectively. Analogously, one can compute the orbital frequency $\Omega\equiv\langle u^{\psi}/u^{v}\rangle$ and the redshift invariant $U\equiv\langle u^{v}\rangle$ [91, 92, 93, 14], [^5] For circular equatorial orbits, these quantities coincide with $\langle u^{\phi}/u^{t}\rangle$ and $\langle u^{t}\rangle$ in Boyer–Lindquist coordinates, since $(v,\psi)$ differ from $(t,\phi)$ only by functions of the radial coordinate $r$ (cfr. Eq. (2.2)).
@@
By solving Eqs. (6.1) at zeroth order in the tidal parameter $\eta$, one recovers the well-known expressions for the radial coordinate, the specific energy, and the specific angular momentum of the unperturbed ISCOs of a Kerr black hole, in agreement with the seminal results of Refs. [5, 75]:
@@
These expressions reveal a strong dependence of the tidal ISCO shifts on both the black-hole spin and the relative orientation of the orbit. In Fig. 1 we illustrate the dependence of these quantities on the spin of the Kerr black hole. We comment further on these results below. In the non-spinning limit $a=0$, all tidal corrections smoothly reduce to those computed for a tidally deformed Schwarzschild black hole in Ref. [73].